# Existence of variational solutions to doubly nonlinear systems in nondecreasing domains

### v1

@@ 1. Introduction
[…]
with parameters $q\in(0,\infty)$ and $p\in(1,\infty)$. For the special choice of $q=1$ and $p=2$, this reduces to the heat equation. In the case of fixed $q=1$ and arbitrary $p\in(1,\infty)$ we are dealing with the parabolic $p$-Laplace equation. Vice versa, for fixed $p=2$ and arbitrary $q\in(0,\infty)$, (1.1) is the porous medium equation (also called fast diffusion equation for $q>1$). More generally, for $p-1>q$, (1.1) is called slow diffusion equation, and for $p-1<q$ it is called fast diffusion equation [31]. This is represented in the behaviour of solutions. Slow diffusion equations allow compactly supported solutions and perturbations propagate with finite speed, while for fast diffusion equations, perturbations propagate with infinite speed, which prevents solutions with compact support.
@@
Here, $b$ is the gradient of a convex $C^{1}$-function satisfying $b(0)=0$ and $a(z,\xi)$ is a continuous and elliptic function in $(z,\xi)$ satisfying a $(p-1)$-growth condition in the second variable. The authors split the lateral boundary $\Omega_{T}$ into two complementary parts, where the first satisfies a Neumann condition and the second a Dirichlet condition. Furthermore, in their proofs the time derivative $\partial_{t}b(u)$ was replaced by the difference quotient $\Delta_{-h}b(u)$, where $h$ is fixed, and a Galerkin method was applied to the elliptic problems. The results in [3] and [30] were generalized to higher order doubly nonlinear equations on unbounded domains by Bernis in [8].
@@
Our goal is to show the existence of variational solutions in noncylindrical domains. We will combine the techniques and methods used in [15] and [17]. However, we weakened certain assumptions and generalized some aspects of the problem. In contrast to [15] the solutions are not restricted to be non-negative. Furthermore, instead of the space-time cylinder $\Omega_{T}$ we will consider a noncylindrical domain, similar to [17]. While the assumptions on $f$ are analogous, cf. (2.4), our primary contribution lies in the analysis of a generalized version of Equation (1.2), where $\partial_{t}u$ is replaced by the nonlinear term $\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}$, which is a special case of $\partial_{t}b(u)$ in [15]. Moreover, unlike the work of Bögelein, Duzaar, Scheven and Singer, which dealt with zero lateral boundary conditions in [17], we impose time-independent lateral boundary values, similar to [15]. Nevertheless, we expect that our methods can be used for time-dependent boundary values. We note that in the case of the doubly nonlinear systems considered here, the question of uniqueness is much more delicate.
@@ 2.1. Setting
The assumptions (2.4)1 and (2.4)2 imply the following lemma, whose proof is analogous to [52, Lemma 2.3].
@@
[…]
@@
$\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}:=|u|^{q-1}u$
@@
$\mathfrak{b}[u,v]$ $:=\tfrac{1}{q+1}|v|^{q+1}-\tfrac{1}{q+1}|u|^{q+1}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot(v-u)$
$=\tfrac{1}{q+1}|v|^{q+1}+\tfrac{q}{q+1}|u|^{q+1}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot v.$
@@
(2.8) $\leq\iint_{E\cap\Omega_{\tau}}f(x,v,Dv)\,\mathrm{d}x\mathrm{d}t+\iint_{E\cap\Omega_{\tau}}\partial_{t}v\cdot(\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}})\,\mathrm{d}x\mathrm{d}t$
@@ 2.2. Main results
in the sense of Definition 2.2. Moreover, we have that $\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\in L^{2}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ and $\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q{+}1$}}}}\in L^{1}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ for any $\varepsilon>0$.
@@
The argument in Section 4.5 shows that $\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\in L^{2}(\Omega_{T},\mathds{R}^{N})$ and $\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q{+}1$}}}}\in L^{1}(\Omega_{T},\mathds{R}^{N})$, provided that in addition to the assumptions of Theorem 2.3 we suppose that
@@
equipped with the norms $\|v\|_{V^{p}(E)}$ and $\|v\|_{V^{p}_{q}(E)}$, respectively. Under some additional conditions on $E$ we show that $\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}$ possesses a distributional time derivative in the dual space $(V^{p,0}(E))^{\prime}$. Indeed, we assume that for all $t\in[0,T)$, the complement $\Omega\setminus E^{t}$ satisfies a measure density condition, i.e., that there exists a constant $\delta>0$ such that there holds
@@
for a.e. $x\in\Omega$ and all $(u,\xi)\in\mathds{R}^{N}\times\mathds{R}^{Nn}$, where $0<\nu\leq L$ and $G\in L^{1}(\Omega_{T})$. Note that (2.10) is a special case of (2.5). In particular, under (2.10), the condition (2.6)3 holds for any $u_{\ast}\in W^{1,p}(\Omega,\mathds{R}^{N})$. Moreover, adapting the proof of [44, Lemma 2.1], we find that (2.4) and (2.10) imply the local Lipschitz condition
@@
Assume that $E$ is a relatively open noncylindrical domain satisfying (2.2) and (2.9), that the variational integrand $f\colon\Omega\times\mathds{R}^{N}\times\mathds{R}^{Nn}\to[0,\infty)$ fulfills (2.4) and  (2.10), and that $u_{\ast}\in W^{1,p}(\Omega,\mathds{R}^{N})\cap L^{q+1}(\Omega,\mathds{R}^{N})$ and $u_{o}\in L^{q+1}(E^{0},\mathds{R}^{N})$. Further, let $u$ be a variational solution to (2.3) in the sense of Definition 2.2. Then, $\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}$ admits a distributional time derivative
$\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\in\big{(}V^{p,0}(E)\big{)}^{\prime}$
@@
$\|\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\|_{(V^{p,0}(E))^{\prime}}\leq c(p,L)\Big{[}\|u\|_{V^{p}(E)}^{p}+\|G\|_{L^{1}(E)}\Big{]}^{\frac{p-1}{p}}.$
@@ 2.3. Methods of proof
Finally, we give the proof of Theorem 2.5 in Section 5 following the approach in [17, Section 4.2.3]. We show that any variational solution is a so-called parabolic minimizer and then use the minimality condition to establish the existence of the time derivative $\partial_{t}(|u|^{q-1}u)$ in the dual space.
@@ 3.1. Technical lemmas
We start this section by deriving estimates that allow us to deal with powers of $u$ and the boundary term $\mathfrak{b}[u,v]$. For the following lemma, we refer to [29, Lemma 8.3]; see also [11, Lemma 3.2] for the statement.
@@
$\tfrac{1}{c(\alpha)}|\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$\alpha$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\alpha$}}}}|\leq(|u|+|v|)^{\alpha-1}|v-u|\leq c(\alpha)|\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$\alpha$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\alpha$}}}}|.$
As a consequence of the preceding lemma, we obtain the following inequality; see also [11, Lemma 3.3].
@@
$|v-u|^{\alpha}\leq c(\alpha)|\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$\alpha$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\alpha$}}}}|.$
@@
$|u-v|^{q+1}\leq c(q)\Big{(}|u|^{\frac{q+1}{2}}+|v|^{\frac{q+1}{2}}\Big{)}\Big{|}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}.$
@@
$=c(q)\Big{|}|u|^{\frac{q+1}{2}}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-|v|^{\frac{q+1}{2}}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}+|v|^{\frac{q+1}{2}}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-|v|^{\frac{q+1}{2}}\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}$
$\leq c(q)\Big{(}|u|^{\frac{q+1}{2}}\Big{|}|u|^{\frac{q+1}{2}}-|v|^{\frac{q+1}{2}}\Big{|}+|v|^{\frac{q+1}{2}}\Big{|}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}\Big{)}$
$\leq c(q)\Big{(}|u|^{\frac{q+1}{2}}+|v|^{\frac{q+1}{2}}\Big{)}\Big{|}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}.\qed$
@@
$\Big{|}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}^{2}\leq c(q)\mathfrak{b}[u,v]\leq c(q)\big{(}\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\cdot(v-u).$
@@
For the first inequality, we refer to [11, Lemma 3.4] with $p$ replaced by $q+1$. Since $w\mapsto\frac{1}{q+1}|w|^{q+1}\in C^{1}(\mathds{R}^{N})$ is convex with $\frac{1}{q+1}D|w|^{q+1}=\boldsymbol{w^{\mbox{\unboldmath{\scriptsize$q$}}}}$, we have that
$\tfrac{1}{q+1}|v|^{q+1}-\tfrac{1}{q+1}|u|^{q+1}\leq\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot(v-u).$
@@
For the proof of the subsequent estimates, cf. [49, Lemma 2.1]; see also [15, Lemma 2.4] for the special case $N=1$.
@@
Let $(u_{k})_{k\in\mathds{N}}\subset L^{q+1}(\Omega,\mathds{R}^{N})$ be a sequence and $u_{\infty}\in L^{q+1}(\Omega,\mathds{R}^{N})$, such that $\boldsymbol{u}_{\boldsymbol{k}}^{q+1}\to\boldsymbol{u}_{\boldsymbol{\infty}}^{q+1}$ in $L^{1}(\Omega,\mathds{R}^{N})$, for $k\to\infty$. Then, we have that $u_{k}\to u_{\infty}$ in $L^{q+1}(\Omega,\mathds{R}^{N})$. Moreover, if $u_{k}\to u_{\infty}$ in $L^{q+1}(\Omega,\mathds{R}^{N})$, for $k\to\infty$, then we have the convergence $\boldsymbol{u_{k}^{\mbox{\unboldmath{\scriptsize$q$}}}}\to\boldsymbol{u_{\infty}^{\mbox{\unboldmath{\scriptsize$q$}}}}$ in $L^{\frac{q+1}{q}}(\Omega,\mathds{R}^{N})$.
@@
For the case $q>1$, we refer to [49, Lemma 2.4]. Thus, it remains to consider the case $10<q<$ 1. The first part of the lemma can be deduced by applying Lemma  3.2 with $\alpha=q+1$, which yields
@@
$\int_{\Omega}\big{|}\boldsymbol{u_{k}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{\infty}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{|}^{\frac{q+1}{q}}\,\mathrm{d}x$ $\leq c(q)\int_{\Omega}\big{|}(\boldsymbol{u_{k}^{\mbox{\unboldmath{\scriptsize$q$}}}})^{\frac{1}{q}}-(\boldsymbol{u_{\infty}^{\mbox{\unboldmath{\scriptsize$q$}}}})^{\frac{1}{q}}\big{|}^{q+1}\,\mathrm{d}x$
@@
We proceed with a Hardy inequality, which holds under a weaker assumption than the measure density condition (2.9). More precisely, denoting the variational $p$-capacity as defined in [34, Def. 5.32] by $\operatorname{cap}_{p}$, we rely on the following concept.
@@
The measure density condition (2.9) with a constant $\delta>0$ implies that $\mathds{R}^{n}\setminus E^{t}$ is uniformly $p$-fat with a parameter $\alpha=\alpha(n,p,\delta)>0$, see, e.g. [34, Example 6.18]. Further, under the measure density condition we have that $\mathcal{V}^{p,0}_{q}(E)=V^{p,0}_{q}(E)$, cf. [17, Lemma 3.1].
For the proof of the following lemma we refer to [35] or [34, Theorem 6.25].
@@
and state the following integration by parts formula. We refer to [49, Lemma 2.11], see also [15, Lemma 2.10] for the case $N=1$ and non-negative solutions.
@@
$\iint_{\Omega_{T}}\Delta_{-h}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot(v-u)\,\mathrm{d}x\mathrm{d}t\leq$ $\iint_{\Omega_{T}}\Delta_{h}v\cdot(\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}})\,\mathrm{d}x\mathrm{d}t$
@@
$\delta_{2}(h)$ $:=\iint_{\Omega\times(-h,0)}\Delta_{h}v\cdot(\boldsymbol{v^{\mbox{\unboldmath{\scriptsize$q$}}}}(t+h)-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}(t))\,\mathrm{d}x\mathrm{d}t.$
@@
We conclude this section by the following compactness result for signed or vector-valued functions. It has been proven in [15, Proposition 3.1] in the case of non-negative functions and in [49, Lemma 2.12] for vector-valued or signed functions.
@@
$\tfrac{1}{h_{\ell}}\sum_{i=1}^{\ell}\int_{\Omega}\Big{|}\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}^{2}\,\mathrm{d}x\leq C$

### v2

@@ 1. Introduction
(1.1) $\partial_{t}(|u|^{q-1}u)-\operatorname{div}(|Du|^{p-2}Du)=0$ $\text{ in }E$
with parameters $q\in(0,\infty)$ and $p\in(1,\infty)$. For the special choice of $q=1$ and $p=2$, this reduces to the heat equation. In the case of fixed $q=1$ and arbitrary $p\in(1,\infty)$ we are dealing with the parabolic $p$-Laplace equation. Vice versa, for fixed $p=2$ and arbitrary $q\in(0,\infty)$, (1.1) is the porous medium equation (also called fast diffusion equation for $q>1$). More generally, for p1>qp-1>q, (1.1) is called slow diffusion equation, and for p1<qp-1<q it is called fast diffusion equation [33]. This is represented in the behaviour of solutions. Slow diffusion equations allow compactly supported solutions and perturbations propagate with finite speed, while for fast diffusion equations, perturbations propagate with infinite speed, which prevents solutions with compact support.
@@
Here, $b$ is the gradient of a convex $C^{1}$-function satisfying $b(0)=0$ and $a(z,\xi)$ is a continuous and elliptic function in $(z,\xi)$ satisfying a (p1)(p-1)-growth condition in the second variable. The authors split the lateral boundary $\Omega_{T}$ into two complementary parts, where the first satisfies a Neumann condition and the second a Dirichlet condition. Furthermore, in their proofs the time derivative $\partial_{t}b(u)$ was replaced by the difference quotient $\Delta_{-h}b(u)$, where $h$ is fixed, and a Galerkin method was applied to the elliptic problems. The results in [4] and [32] were generalized to higher order doubly nonlinear equations on unbounded domains by Bernis in [9].
@@
Our goal is to show the existence of variational solutions in noncylindrical domains. We will combine the techniques and methods used in [17] and [19]. However, we weakened certain assumptions and generalized some aspects of the problem. In contrast to [17] the solutions are not restricted to be non-negative. Furthermore, instead of the space-time cylinder $\Omega_{T}$ we will consider a noncylindrical domain, similar to [19]. While the assumptions on $f$ are analogous, cf. (2.4), our primary contribution lies in the analysis of a generalized version of Equation (1.2), where $\partial_{t}u$ is replaced by the nonlinear term $\partial_{t}\big(|u|^{q-1}u\big)$, which is a special case of $\partial_{t}b(u)$ in [17]. Moreover, unlike the work of Bögelein, Duzaar, Scheven and Singer, which dealt with zero lateral boundary conditions in [19], we impose time-independent lateral boundary values, similar to [17]. Nevertheless, we expect that our methods can be used for time-dependent boundary values. We note that in the case of the doubly nonlinear systems considered here, the question of uniqueness is much more delicate. In particular, the inclusion of coefficients is already difficult in the case of cylindrical domains. However, for filtration equations that include the prototype porous medium equation as a special case, uniqueness can be proven by testing with a Oleĭnik type test function, see [60, Theorem 5.13]. In noncylindrical domains, Abdulla [1] established uniqueness for non-negative solutions to the prototype porous medium equation with non-negative continuous boundary values. If $E$ is a finite union of cylinders, the assumptions on the regularity of the lateral boundary were weakened by Björn, Björn, Gianazza and Siljander [12]. Moreover, doubly nonlinear equations that are independent of $(x,t)$ were investigated by Otto [49] in the cylindrical setting. He proved uniqueness of solutions to the Dirichlet problem with time independent boundary values via an $L^{1}$-contraction principle.
@@ 2.1. Setting
We point out that we do not require a growth assumption from above. However, the assumptions (2.4)1 and (2.4)2 imply the following result, which can be seen as a mild growth property. Its proof is analogous to [55, Lemma 2.3]. It shows that the integrability in $x$ is uniform with respect to $(u,\xi)$ in any bounded set in $\mathds{R}^{N}\times\mathds{R}^{Nn}$.
@@
(2.6) $\left\{\begin{array}[]{l}\mbox{$u_{o}\in L^{q+1}(\Omega,\mathds{R}^{N}$),}\\[5.0pt] \mbox{$u_{\ast}\in L^{q+1}(\Omega,\mathds{R}^{N})\cap W^{1,p}(\Omega,\mathds{R}^{N})$,}\\[5.0pt] \mbox{$0\leq\int_{\Omega}f(x,u_{\ast}+\phi,D(u_{\ast}+\phi))\,\mathrm{d}x<\infty,\quad\forall\phi\in C_{0}^{\infty}(\Omega,\mathds{R}^{N})$.}\end{array}\right$.
@@
$\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}:=|u|^{q-1}u$
@@
$\mathfrak{b}[u,v]$ $:=\tfrac{1}{q+1}|v|^{q+1}-\tfrac{1}{q+1}|u|^{q+1}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\cdot(v-u)$
$=\tfrac{1}{q+1}|v|^{q+1}+\tfrac{q}{q+1}|u|^{q+1}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\cdot v.$
@@
(2.8) $\leq\iint_{E\cap\Omega_{\tau}}f(x,v,Dv)\,\mathrm{d}x\mathrm{d}t+\iint_{E\cap\Omega_{\tau}}\partial_{t}v\cdot(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}})\,\mathrm{d}x\mathrm{d}t$
@@ 2.2. Main results
in the sense of Definition 2.2. Moreover, we have that $\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\in L^{2}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ and $\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$q{+}1$}}}}\in L^{1}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ for any $\varepsilon>0$.
@@
The argument in Section 4.5 shows that $\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\in L^{2}(\Omega_{T},\mathds{R}^{N})$ and $\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$q{+}1$}}}}\in L^{1}(\Omega_{T},\mathds{R}^{N})$, provided that in addition to the assumptions of Theorem 2.3 we suppose that
@@
equipped with the norms $\|v\|_{V^{p}(E)}$ and $\|v\|_{V^{p}_{q}(E)}$, respectively. Under some additional conditions on $E$ we show that $\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}$ possesses a distributional time derivative in the dual space $(V^{p,0}(E))^{\prime}$. Indeed, we assume that for all $t\in[0,T)$, the complement $\Omega\setminus E^{t}$ satisfies a measure density condition, i.e., that there exists a constant $\delta>0$ such that there holds
@@
for a.e. $x\in\Omega$ and all $(u,\xi)\in\mathds{R}^{N}\times\mathds{R}^{Nn}$, where $0<\nu\leq L$ and $G\in L^{1}(\Omega_{T})$. Note that (2.10) is a special case of (2.5). In particular, under (2.10), the condition (2.6)3 holds for any $u_{\ast}\in W^{1,p}(\Omega,\mathds{R}^{N})$. Moreover, adapting the proof of [46, Lemma 2.1], we find that (2.4) and (2.10) imply the local Lipschitz condition
@@
Assume that $E$ is a relatively open noncylindrical domain satisfying (2.2) and (2.9), that the variational integrand $f\colon\Omega\times\mathds{R}^{N}\times\mathds{R}^{Nn}\to[0,\infty)$ fulfills (2.4) and  (2.10), and that $u_{\ast}\in W^{1,p}(\Omega,\mathds{R}^{N})\cap L^{q+1}(\Omega,\mathds{R}^{N})$ and $u_{o}\in L^{q+1}(E^{0},\mathds{R}^{N})$. Further, let $u$ be a variational solution to (2.3) in the sense of Definition 2.2. Then, $\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}$ admits a distributional time derivative
$\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\in\big(V^{p,0}(E)\big)^{\prime}$
@@
$\|\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\|_{(V^{p,0}(E))^{\prime}}\leq c(p,L)\Big[\|u\|_{V^{p}(E)}^{p}+\|G\|_{L^{1}(\Omega)}\Big]^{\frac{p-1}{p}}.$
@@ 2.3. Methods of proof
Finally, we give the proof of Theorem 2.5 in Section 5 following the approach in [19, Section 4.2.3]. We show that any variational solution is a so-called parabolic minimizer and then use the minimality condition to establish the existence of the time derivative $\partial_{t}(|u|^{q-1}u)$ in the dual space.
@@ 3.1. Technical lemmas
We start this section by deriving estimates that allow us to deal with powers of $u$ and the boundary term $\mathfrak{b}[u,v]$. For the following lemma, we refer to [31, Lemma 8.3]; see also [13, Lemma 3.2] for the statement.
@@
$\tfrac{1}{c(\alpha)}|\boldsymbol{v^{\mbox{{\scriptsize$\alpha$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$\alpha$}}}}|\leq(|u|+|v|)^{\alpha-1}|v-u|\leq c(\alpha)|\boldsymbol{v^{\mbox{{\scriptsize$\alpha$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$\alpha$}}}}|.$
As a consequence of the preceding lemma, we obtain the following inequality; see also [13, Lemma 3.3].
@@
$|v-u|^{\alpha}\leq c(\alpha)|\boldsymbol{v^{\mbox{{\scriptsize$\alpha$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$\alpha$}}}}|.$
@@
$|u-v|^{q+1}\leq c(q)\Big(|u|^{\frac{q+1}{2}}+|v|^{\frac{q+1}{2}}\Big)\Big|\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{v^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|.$
@@
$=c(q)\Big||u|^{\frac{q+1}{2}}\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-|v|^{\frac{q+1}{2}}\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}+|v|^{\frac{q+1}{2}}\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-|v|^{\frac{q+1}{2}}\boldsymbol{v^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|$
$\leq c(q)\Big(|u|^{\frac{q+1}{2}}\Big||u|^{\frac{q+1}{2}}-|v|^{\frac{q+1}{2}}\Big|+|v|^{\frac{q+1}{2}}\Big|\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{v^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|\Big)$
$\leq c(q)\Big(|u|^{\frac{q+1}{2}}+|v|^{\frac{q+1}{2}}\Big)\Big|\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{v^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|.\qed$
@@
$\Big|\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{v^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|^{2}\leq c(q)\mathfrak{b}[u,v]\leq c(q)\big(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\cdot(v-u).$
@@
For the first inequality, we refer to [13, Lemma 3.4] with $p$ replaced by $q+1$. Since $w\mapsto\frac{1}{q+1}|w|^{q+1}\in C^{1}(\mathds{R}^{N})$ is convex with $\frac{1}{q+1}D|w|^{q+1}=\boldsymbol{w^{\mbox{{\scriptsize$q$}}}}$, we have that
$\tfrac{1}{q+1}|v|^{q+1}-\tfrac{1}{q+1}|u|^{q+1}\leq\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}\cdot(v-u).$
@@
For the proof of the subsequent estimates, cf. [52, Lemma 2.1]; see also [17, Lemma 2.4] for the special case $N=1$.
@@
Let $(u_{k})_{k\in\mathds{N}}\subset L^{q+1}(\Omega,\mathds{R}^{N})$ be a sequence and $u_{\infty}\in L^{q+1}(\Omega,\mathds{R}^{N})$, such that $\boldsymbol{u}_{\boldsymbol{k}}^{q+1}\to\boldsymbol{u}_{\boldsymbol{\infty}}^{q+1}$ in $L^{1}(\Omega,\mathds{R}^{N})$, for $k\to\infty$. Then, we have that $u_{k}\to u_{\infty}$ in $L^{q+1}(\Omega,\mathds{R}^{N})$. Moreover, if $u_{k}\to u_{\infty}$ in $L^{q+1}(\Omega,\mathds{R}^{N})$, for $k\to\infty$, then we have the convergence $\boldsymbol{u_{k}^{\mbox{{\scriptsize$q$}}}}\to\boldsymbol{u_{\infty}^{\mbox{{\scriptsize$q$}}}}$ in $L^{\frac{q+1}{q}}(\Omega,\mathds{R}^{N})$.
@@
For the case $q>1$, we refer to [52, Lemma 2.4]. Thus, it remains to consider the case $0<q<1$. The first part of the lemma can be deduced by applying Lemma  3.2 with $\alpha=q+1$, which yields
@@
$\int_{\Omega}\big|\boldsymbol{u_{k}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\infty}^{\mbox{{\scriptsize$q$}}}}\big|^{\frac{q+1}{q}}\,\mathrm{d}x$ $\leq c(q)\int_{\Omega}\big|(\boldsymbol{u_{k}^{\mbox{{\scriptsize$q$}}}})^{\frac{1}{q}}-(\boldsymbol{u_{\infty}^{\mbox{{\scriptsize$q$}}}})^{\frac{1}{q}}\big|^{q+1}\,\mathrm{d}x$
@@
We proceed with a Hardy inequality, which holds under a weaker assumption than the measure density condition (2.9). More precisely, denoting the variational $p$-capacity as defined in [36, Def. 5.32] by $\operatorname{cap}_{p}$, we rely on the following concept.
@@
The measure density condition (2.9) with a constant $\delta>0$ implies that $\mathds{R}^{n}\setminus E^{t}$ is uniformly $p$-fat with a parameter $\alpha=\alpha(n,p,\delta)>0$, see, e.g. [36, Example 6.18]. Further, under the measure density condition we have that $\mathcal{V}^{p,0}_{q}(E)=V^{p,0}_{q}(E)$, cf. [19, Lemma 3.1].
For the proof of the following lemma we refer to [37] or [36, Theorem 6.25].
@@
and state the following integration by parts formula. We refer to [52, Lemma 2.11], see also [17, Lemma 2.10] for the case $N=1$ and non-negative solutions.
@@
$\iint_{\Omega_{T}}\Delta_{-h}\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\cdot(v-u)\,\mathrm{d}x\mathrm{d}t\leq$ $\iint_{\Omega_{T}}\Delta_{h}v\cdot(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}})\,\mathrm{d}x\mathrm{d}t$
@@
$\delta_{2}(h)$ $:=\iint_{\Omega\times(-h,0)}\Delta_{h}v\cdot(\boldsymbol{v^{\mbox{{\scriptsize$q$}}}}(t+h)-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}(t))\,\mathrm{d}x\mathrm{d}t.$
@@
We conclude this section by the following compactness result for signed or vector-valued functions. It has been proven in [17, Proposition 3.1] in the case of non-negative functions and in [52, Lemma 2.12] for vector-valued or signed functions.
@@
$\tfrac{1}{h_{\ell}}\sum_{i=1}^{\ell}\int_{\Omega}\Big|\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|^{2}\,\mathrm{d}x\leq C$